# Uncertainty-enabled machine learning for emulation of regional sea-level change caused by the Antarctic Ice Sheet

### v1

@@ 4 Uncertainty quantification
where $\hat{y}_{ij}^{test}$ denotes the prediction from the ML model for the $i^{th}$ location and $j^{th}$ test sample, $i=1,..,27$, $j=1,...,119$. Additionally, $\bar{\hat{y}}_{i}^{val}$ and $\widetilde{y}_{ij}^{test}$ indicate the location-wise mean of the prediction from the ML model $\hat{y}_{ij}^{val}$ and the calibrated point estimate derived through a simple linear regression:
@@
where $y_{ij}^{val}$ represents the (simulated) sea-level change for the $i^{th}$ location and $j^{th}$ validation sample, while $\widetilde{y}_{ij}^{val}$ denotes the calibrated point estimate obtained through linear regression. Lastly, $t_{1-\alpha/2,n-2}$ denotes the 1-$\alpha$ quantile of a $t$ distribution with $n-2$ degree of freedom, where $n=119$ represents the sample size (i.e., the validation set size). We emphasize that we utilize the validation set not only to monitor potential overfitting but also to derive estimates for a simple linear regression. In other words, $\hat{y}_{ij}^{val}$, $\bar{\hat{y}}_{i}^{val}$, $S_{i}^{2,val}$, $\hat{\beta}_{i,0}^{val}$, and $\hat{\beta}_{i,1}^{val}$ are estimated based on the validation set, and these estimates are then employed to calibrate the ML predictions.
@@
So long as a linear model is an appropriate fit between observations (i.e., the sea-level change simulations) and ML predictions, prediction intervals should cover with nominal probabilities. In Section 5 that follows, we demonstrate empirical coverage rates on a hold-out test set of simulations. While we do not explore alternate residual distributions or spatial correlation in the residuals, we believe these items would be worth exploring in future works.
@@ 5 Emulator results
Figure 3: R-squared comparison between calibrated ML model predictions and actual simulated sea-level change values. R-squared values indicate comparable accuracy for the NN, CVAE, and GP models, though the RF model performs worst.
Figure 4: Comparison of empirical prediction interval coverage rates, suggesting that the CVAE and NN have coverage closest to the nominal rate. Note that since there are on the order of 100 examples in the test hold-out set, it is difficult to distinguish between coverage rates of 1%.
Our main results consist of a comparison of accuracy, empirical prediction interval coverage rate, and prediction interval lengths computed on the test hold-out set. We calculate Pearson’s correlation coefficient of determination, $R^{2}$, between calibrated ML model predictions and the actual simulated sea-level change values as a measure of accuracy. The boxplots of Figure 3, 4, and 5 depict the range of values for each of the models across the 27 distinct global locations.
From Figure 3, it is evident that the CVAE, GP, and NN models exhibit similar prediction performance, while the RF model shows inferior performance. The average $R^{2}$ values for each model are 0.972 (NN), 0.969 (CVAE), 0.971 (GP), and 0.844 (RF). This suggests that the CVAE and NN can serve as alternatives to GP-based emulators, which have been widely used.
Figure 5: Average length comparison of prediction intervals in meters – the Gaussian process intervals are shortest but are overconfident as based on coverage rate. The RF tends to be longest.
Figure 4 and 5 show the range of (1-$\alpha$)empirical coverage rates and the length of prediction intervals for the significance level $\alpha$=0.01, 0.05, 0.1, and 0.15. From Figure 4, one can see that the empirical coverage rates for the GP model underestimate the nominal rate more than other ML methods after calibration. It is noteworthy that the RF still exhibits some deviation in empirical coverage from the nominal rates, even after calibration. This is due to the less accurate prediction performance of the RF model. While the aim of calibration is to enhance the accuracy of prediction intervals and point estimates, its effectiveness can be compromised if the underlying model lacks prediction performance.
Figure 5 indicates that the lengths of the GP prediction intervals are generally the shortest but compensate with undercoverage; i.e., a smaller empirical coverage probability than is nominally stated for the four levels shown. On the other hand, due to the overall inaccuracy, the RF tends to have the longest intervals as calculated with simple linear regression, emphasizing the importance of the underlying model.
The prediction intervals from the NN and CVAE are overall quite similar in terms of length, coverage, and accuracy. It appears that the CVAE is essentially behaving almost identically to the NN; we believe this is because the conditional ice-thickness input is providing much of the information in comparison to the latent variable for the prediction, and the decoder architecture of the CVAE is the same as the NN. Overall, the NN and CVAE models seem to perform best taking into account all three characteristics being compared.
As an illustration of the simple linear regression approach for determining prediction intervals, Figure 6 shows a scatterplot depicting simulated sea-level change on the $y$-axis and NN predictions on the $x$-axis for the location Balboa. It is noticeable that the $y=x$ line (solid red) closely aligns with the simple regression line (solid blue), indicating a high degree of accuracy in the predictions ($R^{2}\approx 0.974$). Furthermore, one can see that a simple linear regression technique can yield well-calibrated 95% prediction intervals (solid green).
Figure 6: A scatter plot featuring predictions on the X-axis and simulated sea-level change on the Y-axis. Solid red line is the $y=x$ line while solid blue is the simple regression line. Solid green lines show prediction intervals for significance level $\alpha=0.05$.
Figure 7: Runtime Comparison. For comparison the full numerical model requires about 60 seconds per solution.
In addition, we considered a variation where the inputs are mean ice-thickness values by distinct geographic sectors in Antarctica where the sectors are illustrated in Figure 9 of Appendix A; using sector-based inputs has a precedent in Van Katwyk et al., (2023). However, for this particular emulator, we found that the sectors performed markedly worse for all models using the sector mean values as input. These comparisons can be seen in the Supplementary Materials.
@@
We additionally assess the runtime for generating predictions with each of the ML models by performing a repeated single prediction task 1,000 times. Note we consider that a single prediction is complete when the model provides the predicted values for all 27 locations. For the CVAE model, the runtime is measured based on a single sample drawn from the assumed latent distribution. The runtime for the GP model is evaluated after its kernel parameters are estimated. Figure 7 indicates that all models can generate predictions within one second. The NN, CVAE, and RF models exhibit comparable and more efficient performance than the GP model. Nonetheless, it is noteworthy that when 500 samples from the latent distribution are utilized for the CVAE, as per our analysis, the runtime extends to approximately 125 seconds. We also observe that the additional calibration step does not contribute to a significant increase in runtime ($\approx$ 0.0006 seconds). Since a single run of the sea-level solver can take about 60 seconds, the demonstrated speed up is greater than two orders of magnitude, particularly for the NN emulator. Training each emulator takes on the order of a few hours on top of the overheard of  21 hours (17 minutes per a single run, times 71 runs) for running the sea-level fingerprint calculations in serial.
6 Illustration of emulator use for a Monte Carlo ensemble
@@
We demonstrate a potential use of the sea-level change emulator for generating distributions of sea-level change outputs for locations across the globe, in order to show how such an emulator could be used to produce RSL distributions from future Antarctic ice-sheet simulations. The most straightforward Monte Carlo approach to do so is to sample many parameters from their respective distributions, run an ice-sheet simulator under the sampled parameters, and subsequently run the resultant ice-sheet states through the sea-level change emulator in order to generate a distribution of sea-level change values. For our purposes, running an ice-sheet simulator iteratively is prohibitively expensive and so we instead construct a Monte Carlo ensemble of ice-sheet states using the ISMIP6 ensemble of Antarctica simulations.
@@
Our goal in the demonstration is to use the neural-network emulator to produce a distribution of sea-level change values for the 27 long-term tide gauge sites. For this illustration, we focus on the year 2100 and high-emissions scenarios, and cull 48 instances from the ISMIP-6 training data matching these criteria. We performed a principal-components analysis on these 48 instances, retaining the top 38 which explained 99% of the variation in the year 2100, high-emissions simulations. We then bootstrapped the principal-component weights 100 times in order to randomly generate a Monte Carlo ensemble of Antarctic ice-sheet states representative of the year 2100 and high emissions. Alternative dimension reduction techniques could be explored, though a principal-components approach was used as an illustrative method to generate a Monte Carlo ensemble. These 100 randomly generated ice-sheet states were then fed into the neural-network emulator. We illustrate emulated densities of sea-level change using the Monte Carlo simulations in Figure 8. We emphasize that these distributions are not calibrated RSL projections and are meant as a demonstration of the application of the NN emulator.
@@
The three sites chosen were selected as they represent among the widest variation in RSL of the cities sampled (Figure 1). The emulated RSL derived from the synthetic ice-sheet states span a similar range of sea-level change as the GMSL estimates from the ISMIP6 ensemble from which they were derived; Seroussi et al., (2020) report mean sea-level contribution of about 0.025 m with a standard deviation of 0.049 m for the high emissions scenarios. Seroussi et al., (2020) also demonstrate a longer tail of positive GMSL change in the ensemble, which our distributions also show (Figure 8). We consider this consistent given we did not use all ISMIP6 samples, our synthetic ice-sheet states are not guaranteed to have the same statistical properties as the ISMIP6 ensemble, we are considering RSL rather than GMSL, and these results come from the NN emulator.
Figure 8: Densities of emulated RSL change for Dunedin, Montevideo, and Midway during year 2100 with high-emissions scenario. These densities are an example of the application of the NN emulator and do not represent calibrated projections of RSL.
@@
While there are clear differences in the densities of RSL projections for the three locations, they share many similarities, despite being from different parts of the Earth. The mode of each density close to zero is a result of our synthetic ice-sheet states generating many samples with small net mass loss from AIS. However, the tails of the distributions exhibit significant differences. Positive RSL values have lower magnitude for the two Southern Hemisphere locations (Dunedin and Montevideo), consistent with the fingerprint of decreased RSL rise closer to AIS (Figure. 1). This has the effect of compressing the range of the density functions. The normalized RSL changes at southern hemisphere locations is more sensitive to the spatial pattern of AIS mass change (Cederberg et al.,, 2023; Roffman et al.,, 2023), but due to the smaller magnitude of RSL change there, also have smaller standard deviations (Cederberg et al.,, 2023). While more detailed evaluation of these synthetic distributions is not appropriate, this example shows the potential utility of the NN emulator for producing RSL estimates rapidly from very large numbers of ice-sheet simulations.

### v2

@@ 5.1 Linear regression prediction intervals
where $\hat{y}_{ij}^{test}$ denotes the prediction from the emulator model for the $i^{th}$ location and $j^{th}$ test sample, $i=1,..,27$, j=1,,119j=1,...,119. Additionally, $\bar{\hat{y}}_{i}^{val}$ and $\widetilde{y}_{ij}^{test}$ indicate the location-wise mean of the prediction from the emulator model $\hat{y}_{ij}^{val}$ and the calibrated point estimate derived through a simple linear regression:
@@
where $y_{ij}^{val}$ represents the (simulated) sea-level change for the $i^{th}$ location and $j^{th}$ validation sample, while $\widetilde{y}_{ij}^{val}$ denotes the point estimate obtained from linear regression. Lastly, $t_{1-\alpha/2,n-2}$ denotes the 1-$\alpha/2$ quantile of a $t$ distribution with n2n-2 degree of freedom, where $n=119$ represents the sample size (i.e., the validation set size). We emphasize that we utilize the validation set not only to monitor potential overfitting but also to derive estimates for a simple linear regression. In other words, $\hat{y}_{ij}^{val}$, $\bar{\hat{y}}_{i}^{val}$, $S_{i}^{2,val}$, $\hat{\beta}_{i,0}^{val}$, and $\hat{\beta}_{i,1}^{val}$ are estimated based on the validation set, and these estimates are then employed to calibrate the emulator predictions.
@@
So long as a linear model is an appropriate fit between observations (i.e., the sea-level change simulations) and emulator predictions, prediction intervals should cover with nominal probabilities. In Section 6 that follows, we demonstrate empirical coverage rates on a hold-out test set of simulations. While we do not explore alternate residual distributions or spatial correlation in the residuals, we believe these items would be worth exploring in future works. Nonetheless, the split-conformal method of UQ next described easily allows for a departure from normality.
5.2 Split-conformal inference
A prevalent modern technique for quantifying predictive uncertainties from ML models is that of split-conformal inference (lei2018distribution). The essence of this approach is to fit the ML model on a training dataset as usual, but use another dataset (e.g., validation) that was not used for training in order to build an empirical distribution of residuals. Then, when making a prediction for unseen test data, one uses a prediction interval $\left[\hat{f}(\bm{x})-d_{\alpha},\hat{f}(\bm{x})+d_{\alpha}\right]$, where $\hat{f}$ is the fitted ML model, $d_{\alpha}$ is a residual chosen from the empirical residual distribution that leads to a specified probability of coverage (i.e., $1-\alpha$), and $\bm{x}$ is a test example. We specifically implement Algorithm 2 of lei2018distribution in computing prediction intervals in the emulator results that follow. We note that while this method has been developed primarily in the context of ML models, it is applicable to any function being fit including the physics-based sensitivity kernel approach.
This method of UQ has several advantages and disadvantages, which are summarized here. Advantages include not having to retrain the ML model as in resampling approaches (i.e., computational efficiency), no distributional assumption on residuals, and a mathematical guarantee of coverage probability meeting the nominal rate; by nominal rate of coverage, we mean the stated value of $1-\alpha$, e.g., 95% for a 95% prediction interval. However, this guarantee is an average over the inputs, meaning that some input cases can have coverage less than the nominal rate as long as the average meets the nominal rate. Relatedly, the interval lengths do not change for different $\bm{x}$ values, i.e., regions of input space. (By prediction interval length, we mean the upper bound of the interval minus the lower bound of the interval.) However, we account for variability in residuals across different regions in this application by using distinct residual distributions for each location, just as we consider different linear regression models for each location discussed previously. Therefore, the lengths of predictive intervals change for the different locations in our application. We note that while this method easily allows for a departure from normality of residuals, it does not explicitly account for biases in the predictions, as is done with the linear-regression intercept term described previously.
5.3 Latent variable sampling and Monte Carlo dropout
The previous two UQ approaches, split-conformal inference and linear regression post-processing, are quite generic and can be applied to all of the emulation methods (i.e., machine learning and physics-based sensitivity kernel) examined in this study. However we note that the CVAE ML model described in Section 3.2 is a generative ML model and implements Monte Carlo (MC) dropout (gal2016dropout), and hence adds another possible method for UQ. MC dropout is implemented in PyTorch and essentially randomly zeroes out connections between neurons in each layer with some pre-specified probability (we used 0.3). Sampling standard normal latent variables and passing them through the decoder of the CVAE also adds another source of variability in predictions. However, when we implemented this method of UQ, we found that the coverage of predictive intervals that are nominally 99+% coverage (i.e., minimum and maximum of samples) covered the true output only 16% or less of the time because the resultant predictive intervals were far too narrow. We note that if MC dropout was used without latent variable sampling, the variability in predictions would be even smaller. The summary statistics for empirical coverage rates and average interval lengths for nominally 99+% intervals over the 27 locations are shown in Tables 1 and 2.
Min 1st Quartile Median Mean 3rd Quartile Max
0.017 0.092 0.118 0.112 0.130 0.168
Table 1: Summary statistics for the empirical coverage rates over the 27 locations of interest using MC dropout for UQ. The prediction intervals here are derived from the minimum and maximum of predictive samples and should be close to .99 or greater if the MC dropout predictive distributions have well-calibrated uncertainties, since 500 samples were generated with MC dropout.
Min 1st Quartile Median Mean 3rd Quartile Max
0.077 0.084 0.087 0.086 0.089 0.095
Table 2: Summary statistics for the average empirical prediction lengths, in cm, over the 27 locations of interest using MC dropout for UQ. While these lengths are very small, because of small coverage rates it can be inferred the MC dropout predictive distributions are very biased, implying poor UQ.
Because these coverages are far less than the nominal rate while the predictive distributions from MC dropout are very narrow, this implies the predictive distributions from the MC dropout approach are very biased and of poor UQ quality. We disregarded this approach for UQ in the comparisons that follow.
@@ 6 Emulator results
Figure 3: R-squared comparison between calibrated emulator predictions and actual simulated sea-level change values. R-squared values indicate that the kernel is most accurate, followed by the CVAE, and then the NN. All emulators show a high level of accuracy; see following linear plots in addition to these boxplots. The variation these display is over the 27 locations examined in this study.
Figure 4: Root mean squared error (RMSE) comparison between calibrated emulator predictions and actual simulated sea-level change values, in units of cm. RMSE values indicate similar insights as the previous figure, in that the kernel is most accurate, followed by the CVAE, and then the NN. The variation these boxplots display is over the 27 locations examined in this study. The median RMSE values are about 0.3 cm for the kernel emulator, 0.5 cm for the CVAE, and 0.7 cm for the NN emulator.
Figure 5: Comparison of empirical prediction interval coverage rates over the three emulator methods (kernel, CVAE, and NN) and two UQ methods (linear model, subscripted as “lm", and split-conformal inference, subscripted as “conf"). For the 85% and 90% prediction intervals, the empirical coverages meet the nominal rate, excepting slight undercoverage for the kernel split-conformal intervals. For the 95% and 99% intervals, the split-conformal methods come closest to meeting the nominal coverage rate. Horizontal dashed line depicts the nominal coverage rate that ideally all methods are close to. The variation in these boxplots is over the 27 locations examined in this study.
Our main results consist of a comparison of accuracy, empirical prediction interval coverage rate, and prediction interval lengths (defined as the upper bound minus the lower bound of the prediction interval) computed on the test hold-out set. We calculate Pearson’s correlation coefficient of determination, $R^{2}$, in addition to root mean squared error (RMSE) between emulator predictions and the actual simulated sea-level change values as a measure of accuracy, though the illustrations in Figure 8 provide more information on accuracy. The boxplots of Figures 3 and 4 depict the range of values for each of the emulators compared (i.e., NN, CVAE, and the sensitivity kernel method) across the 27 distinct global locations of this study. From Figures 3 and 4, it is evident the sensitivity kernel is most accurate, followed by the CVAE, and finally the NN. However, the CVAE is close to the sensitivity kernel, and for 6 locations achieves a larger $R^{2}$ value. We note that we also considered a variation where the inputs are not flattened vectors but instead based on sectors in Antarctica (see A), though the accuracy in that case was markedly worse than the results here, so we did not pursue it further.
We also performed paired Wilcoxon signed-rank tests to statistically compare accuracy, measured via $R^{2}$ and RMSE, between the kernel and CVAE and kernel and NN emulators. The Wilcoxon signed-rank test is nonparametric and does not assume normality. We paired both $R^{2}$ and RMSE for each of the 27 locations when conducting these statistical tests. The p-values of all of the tests are very small, with the largest being $1.76\times 10^{-5}$, supporting the alternative that the kernel emulator is more accurate (in terms of smaller RMSE and larger $R^{2}$) than either the NN or CVAE emulators.
Figure 6: Comparison of average interval lengths, in meters, comparing the three emulator methods (kernel, CVAE, and NN) and two UQ methods (linear model, subscripted as “lm", and split-conformal inference, subscripted as “conf"). Note that the coverage of the split-conformal method in the 99% case comes at a substantial length tradeoff for the CVAE. The variation in these boxplots is over the 27 locations examined in this study.
Figure 7: Comparison of average interval scores, comparing the three emulator methods (kernel, CVAE, and NN) and two UQ methods (linear model, subscripted as “lm", and split-conformal inference, subscripted as “conf"). While the scores are generally better (lower score is better) for the conformal method than the linear model, this is not apparent for the CVAE and kernel methods for the 99 percent prediction intervals. The longer prediction intervals for the 99 percent conformal case can explain this behavior especially for the CVAE. The variation in these boxplots is over the 27 locations examined in this study.
We compare the empirical coverage rates and prediction interval lengths as a means of assessing the quality of UQ for all three emulator methods and for both the split-conformal and linear-regression methods of UQ. We show those for nominal miscoverage values of $\alpha=0.01,0.05,0.10$, and $0.15$ in Figures 5 and 6. Our findings suggest that the best coverages are achieved by the split-conformal method, though the lengths of the intervals, particularly for $\alpha=0.01$, may be too large to be of practical value in applications. The lengths of predictive intervals for the linear regression method are far more useful (in terms of length) for smaller values of $\alpha$ but these intervals undercover especially for $\alpha=0.01$. In addition to empirical coverage rates and prediction interval lengths, we also compute the average proper interval score (gneiting2007strictly) per location and display these in Figure 7; the findings here are similar to the analysis of coverage and length, in that the split-conformal scores are general lower (i.e., better), but in the 99% case split-conformal and the linear model approach are nearly tied for the kernel and CVAE methods, and slightly larger for the NN method using the linear model approach. As is reviewed more in gneiting2007strictly, the proper interval scoring rule combines empirical coverage and interval length into a single metric, which penalizes prediction interval length as well as empirical miscoverage relative to the nominal rate. For example, if the true value being predicted is a fixed distance away from the lower bound of a prediction interval, this will be penalized more for a higher nominal coverage rate.
Figure 8 shows an illustration depicting the linear-regression results used for UQ for three locations that represent a wide variation of RSL change amongst the 27 cities: Midway, Montevideo, and Dunedin. The three sites chosen were selected as they represent among the widest variation in RSL of the cities sampled (Figure 1). Simulated sea-level change (i.e., the prediction target) is on the $y$-axis and emulator predictions are on the $x$-axis. Each column represents the three emulator methods in this study: NN, CVAE, and the kernel method. Additionally, we plot histograms of residuals derived for use in split-conformal inference in Figure 9. These residuals suggest that the scale of errors is smallest for the kernel method, further corroborating the claim that it is the most accurate of the emulation methods compared.
Figure 8: Scatterplots comparing sea-level change simulation targets (y-axis) with emulator predictions (x-axis) on the test set for Dunedin, Montevideo, and Midway, in meters. Solid line indicates when predictions equal simulations, i.e. a perfect fit. The kernel method is most accurate, followed by the CVAE, and then NN. The kernel method also extrapolates best – it is most accurate for the most extreme point for all three locations.
Figure 9: Empirical distributions of residuals in meters used for split-conformal inference illustrated with histograms at the same three locations: Dunedin, Montevideo, and Midway, which showcase a wide spectrum of RSL variation. Consistent with the previously shown scatterplots, the residuals are smallest in magnitude for the kernel method. The residual distributions are quite similar for the CVAE and NN, despite that the CVAE appears to have substantially higher R-squared values.
@@
We additionally assess the runtime for generating predictions with each of the ML models by performing a single prediction task 1,000 times. All 1,000 NN predictions completed in less than 0.25 seconds and all but one of the CVAE runs completed in less than 0.375 seconds. We consider that a single prediction is complete when the model provides the predicted values for all 27 locations. For the CVAE model, the runtime is measured based on a single sample drawn from the assumed latent distribution. We also observe that the additional calibration step does not contribute to a significant increase in runtime ($\approx$ 0.0006 seconds). Training each of the ML emulators takes on the order of a few hours on top of the overhead of  21 hours (17 minutes per a single run, times 71 runs) for running the sea-level fingerprint calculations in serial. In contrast, the kernel method’s runtime is due to a matrix-vector multiplication and requires only a single forward solve for generating the sensitivity kernel and thus substantially much less overhead than the NN and CVAE for training. The kernel method takes a mean and median of 5.2 seconds to generate emulated sea-level change values for all 27 locations of interest for all 17 time levels; thus taking 0.3 seconds to make a prediction at all 27 locations, which is close to the CVAE and NN runtime. Additionally, the kernel method took 33 seconds to generate the kernels for the 27 cities. Given that the full solve takes about 60 seconds per time level, all emulation methods afford a speed up of more than 100 times over running the full simulation for generating a sea-level fingerprint; we stress, however, that the kernel-generation overhead for the kernel method is substantially less than the training overhead for the ML methods.
In Table 3, we summarize the computational costs for the emulators for three steps: 1) generating the sea-level fingerprints via the numerical sea-level solver, training the emulators on the training data, and evaluating the emulators to make a prediction for the 27 locations of interest.
Emulator Data generation (min) Emulator fitting (min) Time for predictions (s)
Sensitivity kernels 1 NA 0.31
NN 969 121 0.25
CVAE 969 286 0.375
Table 3: Summary of computational runtime comparisons for generating sea-level fingerprints used for building emulators, training the emulators, and making predictions for 27 locations of interest, after the emulators have been trained. Note that for the kernel method, the kernel for each location is created upon running the sea-level solver such that no additional training is needed (in contrast to the ML methods). The ML models were trained using a shared graphics processing unit (GPU) partition available from the Darwin (osti_1441285) cluster at Los Alamos National Laboratory.
7 Application of emulators for a Monte Carlo ensemble
@@
We demonstrate a potential use of the sea-level change emulators for generating distributions of simulated sea-level change outputs for locations across the globe, in order to show how such an emulator could be used to produce RSL distributions from future Antarctic ice-sheet simulations. The most straightforward Monte Carlo approach to do so is to sample many parameters from their respective distributions, run an ice-sheet simulator under the sampled parameters, and subsequently run the resultant ice-sheet states through the sea-level change emulator in order to generate a distribution of sea-level change values. For our purposes, running an ice-sheet simulator iteratively is prohibitively expensive and so we instead construct a Monte Carlo ensemble of ice-sheet states using the ISMIP6 ensemble of Antarctica simulations.
@@
Our goal in the demonstration is to use the kernel and NN emulators to produce a distribution of sea-level change values for the 27 long-term tide gauge sites. For this example, we focus on the year 2100 and high-emissions scenarios, and cull 48 instances from the ISMIP6 training data matching these criteria. We performed a principal-components analysis on these 48 instances, retaining the top 32 which explained 99% of the variation in the year 2100, high-emissions simulations. We then bootstrapped the principal-component weights 100 times in order to randomly generate a Monte Carlo ensemble of Antarctic ice-sheet states representative of the year 2100 and high emissions. Alternative dimension reduction techniques could be explored, though a principal-components approach was used as a practical method to generate a Monte Carlo ensemble. These 100 randomly generated ice-sheet states were then fed into the kernel and neural-network emulators. We illustrate emulated densities of sea-level change using the Monte Carlo simulations in Figures 10 and 11. We emphasize that these distributions are just meant as a demonstration of the application of the emulators.
@@
As in Section 6, the three locations were selected because they are indicative of the widest variation in RSL of the cities sampled (Figure 1). The emulated RSL derived from the synthetic ice-sheet states span a similar range of sea-level change as the GMSL estimates from the ISMIP6 ensemble from which they were derived; Seroussi2020 report mean sea-level contribution of about 0.025 m with a standard deviation of 0.049 m for the high emissions scenarios. Seroussi2020 also demonstrate a longer tail of positive GMSL change in the ensemble, which our distributions also show (Figures 10 and 11). We consider this consistent given we did not use all ISMIP6 samples, our synthetic ice-sheet states are not guaranteed to have the same statistical properties as the ISMIP6 ensemble, we are considering RSL rather than GMSL, and these results come from the kernel and NN emulators.
Figure 10: Probability density plots of emulated RSL change for Dunedin, Montevideo, and Midway during year 2100 with high-emissions scenario, produced with the kernel emulator, which is most accurate of all emulators compared previously. We expect Midway to have the heaviest right tail given its location in the Pacific and Dunedin to have the lightest right tail. Both aspects are evidenced here. Note that these are densities for simulated RSL change and are an example of the application of the emulator. RSL units in meters.
Figure 11: Probability density plots of emulated RSL change for Dunedin, Montevideo, and Midway during year 2100 with high-emissions scenario, produced with the NN emulator, which is least accurate of all emulators compared previously. The right tails of these densities are similar to the previous plots in that Midway has the heaviest right tail whereas Dunedin has the lightest right tail, though the difference is most pronounced in the previous densities – which are more reliable due to the kernel method’s better extrapolation ability. Note that these are densities for simulated RSL change and are an example of the application of the emulator. RSL units are in meters.
@@
While there are clear differences in the densities of RSL projections for the three locations, they share many similarities, despite being from different parts of the Earth. The mode of each density close to zero is a result of our synthetic ice-sheet states generating many samples with small net mass loss from AIS. However, the tails of the distributions exhibit significant differences. Positive RSL values have lower magnitude for the two Southern Hemisphere locations (Dunedin and Montevideo), consistent with the fingerprint of decreased RSL rise closer to AIS (Figure. 1). This has the effect of compressing the range of the density functions. The normalized RSL changes at southern hemisphere locations is more sensitive to the spatial pattern of AIS mass change (Cederberg2023; Roffman2023), but due to the smaller magnitude of RSL change there, also have smaller standard deviations (Cederberg2023). While more detailed evaluation of these synthetic distributions is outside the scope of this work, this example shows the potential utility of emulators for producing RSL estimates rapidly from very large numbers of ice-sheet simulations.